Calculate the ΔG°rxn at 298 K using the following information.

                        2 HNO3(aq) + NO(g) → 3 NO2(g) + H2O(l)           ΔG°rxn = ?

ΔH°f (kJ/mol) -207.0 91.3 33.2   -285.8

S°(J/mol∙K        146.0 210.8 240.1   70.0
Answer #1

2 HNO3(aq) + NO(g) → 3 NO2(g) + H2O(l)    

ΔH°rxn   = ΔH°f products - ΔH°f reactants

              = (-285.8 + 3*33.2) - (2*-207 + 91.3)

               = 136.5 KJ/mole

ΔS°rxn   = S°f products - S°f reactants

               = 70 + 3*240.1 -(2*146+210.8)

                = 287.5J/mole-K   = 0.2875Kj/mole-K

ΔG°rxn     = ΔH°rxn- TΔS°rxn

                = 136.5 - 298*0.2875   = 50.825 KJ/mole

2 HNO3(aq) + NO(g) → 3 NO2(g) + H2O(l)           ΔG°rxn = 50.825 KJ/mole